OpenStudy (anonymous):

Find the value of the real number y if the imaginary part and the real part of (1+3i)/(4-iy) are equal. Please help me solve this problem:) and can u tell me the explanation too pleasee? thankyou soo much!

OpenStudy (anonymous):

i get y = -2

OpenStudy (anonymous):

how do u get the answerr?

OpenStudy (anonymous):

not sure but here is what i did \[\frac{1+3i}{4-yi}=\frac{1+3i}{4-yi}\times \frac{4+yi}{4+yi}=\frac{-3y+4+(y+12)i}{16+y^2}\]

OpenStudy (anonymous):

or \[\frac{-3y+4}{16+y^2}+\frac{y+12}{16+y^2}i\] and since real and imaginary parts are the same you should have \[-3y+4=y+12\]

OpenStudy (anonymous):

thankyou :)) how about \[\left( 1+3y \right) / \left( 4-iy \right)\]

OpenStudy (anonymous):

for that second one:\[y=-\frac{1}{3}\ or\ y=4\]

OpenStudy (anonymous):

can you also post the explanations pleasee? i tried to times (4+iy)/(4+iy) to the question but i cant solve ittt :((

OpenStudy (anonymous):

if you times it you get \[\frac{4(3y+1)}{y^2+16}+\frac{y(3y+1)}{y^2+16}i\] since both real and imaginary parts are the same \[12y+4=3y^2+y\] solve for y

OpenStudy (anonymous):

\[3y^2-11y-4=0\]

OpenStudy (anonymous):

thankyou soo muchh:DD